# Infra-red thermographic inversion in ST40

### v1

@@ Acknowledgments
This work was made possible by the US-DOE CRADA NFE-19-07769. Acknowledgment to the ever-precious help of David and Gabriela Hoffman, Sam Suchal, Teresa Thornton and Rosanne Monteiro is duely given.
@@
The authors have no conflicts to disclose. The authors’ contribution statement using CRediT reads as follows. Matteo Moscheni: conceptualization (equal); methodology (equal); software (lead); formal analysis (lead); data curation (equal); visualization (lead); writing - original draft (lead); writing - review and editing (equal). Erik Maartensson: conceptualization (equal); methodology (equal); software (equal); writing - review and editing (equal). Matthew Robinson: software (equal); formal analysis (equal); writing - review and editing (equal). Chris Marsden: conceptualization (equal); software (equal); writing - review and editing (equal). Adrian Rengle: resources (lead); data curation (equal); writing - review and editing (equal). Travis Kelly Gray: conceptualization (equal); resources (equal); supervision (equal); writing - review and editing (equal). Andrea Scarabosio: supervision (equal); writing - review and editing (equal). Elena Vekshina: supervision (equal); writing - review and editing (equal). Xin Zhang: supervision (equal); writing - review and editing (equal).
@@
Appendix A: Calcam
The temperature information $T_{\text{IR}}$ from the IR camera view (Fig. 1b) belongs to the 2D pixel space, i.e. $T_{\text{IR}}=T_{\text{IR}}(x_{\text{p}},\phi_{\text{p}})$, with $x_{\text{p}}\in[1;540]$ and $\phi_{\text{p}}\in[1;612]$ (Sec. 2.2). During the pre-processing phase, the Calcam software [47] is employed to map such information in the three-dimensional CAD space ($\sim$ real space). The planar sub-space $x\phi$ of the CAD surface of the divertor tiles (Sec. 3.1.2) is ultimately of interest in FAHF. Calcam therefore creates the univocal relationship:
Calcam: $T_{\text{IR}}(x_{\text{p}},\phi_{\text{p}})\mapsto$ CAD space $\mapsto T_{\text{IR}}(x,\phi)$. (13)
@@
Appendix B: Python3 implementation
The entirety of the FAHF algorithms is written in Python3 [60], also because of its versatility when interfacing with third-party softwares (e.g. Calcam [47], Appendix A). Acknowledging the usual lack of implementation-specific information in the literature dealing with IR thermographical inversions, the actual implementation of Eq. (4) in Python3 is here concisely reported.
The meaning of the symbols is obvious, in fact coherent with the mathematical notation adopted throughout the paper. The only exception is for the indeces which are diminshed by 1, because of the Python3 convention of indexing from 0. {python}
@@
Notably, scipy’s function convolve2d is easily parallelised across multiple cores.

### v2

@@ Acknowledgments
The submitted manuscript has been co-authored by a contractor of the U.S. Government under contract DE-AC05-00OR22725. Accordingly, the U.S. Government retains a nonexclusive, royalty-free license to publish or reproduce the published form of this contribution, or allow others to do so, for U.S. Government purposes.
We then thank OpenAI’s language model, ChatGPT, for providing insightful editorial feedback on early drafts of this manuscript.
Last, acknowledgment to the ever-precious help of David and Gabriela Hoffman, Sam Suchal, Teresa Thornton, Rosanne Monteiro and Anne Marie Frohock is also duely given.
@@
The authors have no conflicts to disclose. The authors’ contribution statement using CRediT reads as follows. Matteo Moscheni: conceptualization (equal); methodology (equal); software (lead); formal analysis (lead); data curation (equal); visualization (lead); writing - original draft (lead); writing - review and editing (equal). Erik Maartensson: conceptualization (equal); methodology (equal); software (equal); writing - review and editing (equal). Matthew Robinson: software (equal); formal analysis (equal). Chris Marsden: software (equal). Adrian Rengle: resources (lead); data curation (equal). Andrea Scarabosio: conceptualization (equal); supervision (equal); writing - review and editing (equal). Patrick Bunting: conceptualization (equal); resources (equal); writing - review and editing (equal). Travis Kelly Gray: conceptualization (equal); resources (equal); supervision (equal); writing - review and editing (equal). Elena Vekshina: supervision (equal); writing - review and editing (equal). Xin Zhang: supervision (equal). Salomon Janhunen: methodology (equal); writing - review and editing (equal).
@@
Calcam
The temperature information $T_{\text{IR}}$ from the IR camera view (Fig. 1b) belongs to the 2D pixel space, i.e. $T_{\text{IR}}=T_{\text{IR}}(\phi_{\text{p}},x_{\text{p}})$, with $\phi_{\text{p}}\in[1;612]$ and $x_{\text{p}}\in[1;540]$ (Sec. 2.2). During the pre-processing phase, Calcam [52], a Python package for geometric calibration of cameras and for performing related analysis, is employed to map such information in the three-dimensional CAD space ($\sim$ real space).
The planar sub-space $\phi x$ of the CAD surface of the divertor tiles (Sec. 4.1.2) is ultimately of interest in FAHF. Calcam therefore creates the univocal relationship:
Calcam: $T_{\text{IR}}(\phi_{\text{p}},x_{\text{p}})\mapsto$ CAD space $\mapsto T_{\text{IR}}(\phi,x)$. (16)
@@
Calcam also conveniently allows: (i) to extract useful data related to the camera LoS, FoV and pupil position (Sec. 2.2), among others; (ii) to account for distortion from rectilinear lenses (e.g. barrelling) and fisheye lenses; (iii) to quantify the camera movement.
Experimental uncertainty
This appendix explores the effect of two different flavours of experimental uncertainty of magnitude up to $\eta^{n}=0.02\times\text{max}_{i\in[1;N_{x}]}\{T_{\text{IR},i}^{n}\}$ on the experimental datum $T_{\text{IR},i}^{n}$, fed as an input to FAHF. Specifically:
  • A random noise within ±$\pm\eta^{n}$ is added to $T_{\text{IR},i}^{n}$. The peculiarly smooth behaviour in space of $T_{\text{IR}}$ in Fig. 1c actually suggests that a random noise would be of even lower entity.
  • A constant shift $+\eta^{n}$ is applied to $T_{\text{IR},i}^{n}$, which can be meant as mimicking a calibration/emissivity correction.
In both cases, an average discrepancy of $2\%$ in the resulting surface heat flux profile in space is found at $t_{*}=128$ ms. Crucially, this does not significantly alter its functional shape (Eich function, Sec. 4.1.6), with the fitting parameters in fact changing by $7\%$ at most -- a figure expected to drop further if the iterative fitting procedure of Ref. [48] was employed (Sec. 3.2).
Caveated by the above, a nominal variation of $10\%$ in $\lambda_{q}$ (still acceptable) is found if the magnitude of the constant shift is increased to $8\%$ ($+31$), with an average discrepancy of 5% over the entire heat flux profile.
In conclusion, a reasonable experimental uncertainty on $T_{\text{IR}}$ would not impact any physical conclusions.
Python3 implementation
The entirety of the FAHF algorithms is written in Python3 [67], also because of its versatility when interfacing with third-party softwares (e.g. Calcam [52], Appendix Calcam). A modular approach is selected, to accommodate future improvements to the code.
Acknowledging the usual lack of implementation-specific information in the literature dealing with IR thermographical inversions, the actual implementation of Eq. (8) in Python3 is here concisely reported.
The meaning of the symbols is obvious, in fact coherent with the mathematical notation adopted throughout the paper. The only exception is for the indeces which are diminshed by 1, because of the Python3 convention of indexing from 0.
{python}
@@
Notably, scipy’s function convolve2d used to compute derivatives[71] is easily parallelised across multiple cores, and even in its serial fashion (30 s per sub-chord, Sec. 5.1.1) allows a 50$\times$ speed-up compared to a for-loop based approach.
Details of time and space convergence
This appendix further expands on the details of the convergence procedures of which in Sec. 4.3.2. In the time convergence, the time-step $\Delta t$ is varied in the interval $[4\times 10^{-6};4\times 10^{-4}]$ s. The actual values are detailed in Table 3. Instead, Table 3 reports the values $\Delta x\in[6.8\times 10^{-5};2.3\times 10^{-3}]$ m of the space convergence. Linear interpolation is employed whenever $\Delta x<\Delta x_{\text{IR}}=2.1\times 10^{-4}$ m (Sec. 2.2.1), and all throughout the time convergence, as $\Delta t$ from Eq. (7) always falls short of $\Delta t_{\text{IR}}=1.25\times 10^{-3}$ s. This does not influence the conclusion of the analyses.
Each simulation in the time convergence is run with a fixed spatial discretisation with $\Delta x=5.5\times 10^{-4}$ m. This space-step is small enough such that to not alter the error in time across all the cases, i.e. $o(\Delta x^{2})\ll o(\Delta t)\;\forall\Delta t$ in the time convergence. In the space convergence, $\Delta t=7.2\times 10^{-6}$ s dually guarantees $o(\Delta t)\ll o(\Delta x^{2})\;\forall\Delta x$.
Common to both procedures, a reference point in time $t_{*}=128$ ms and a reference point in space $(x_{*},y_{*})=(16.55;2.21)$ mm are selected where to locally compare the temperature across the various discretisations. The time point $t_{*}=128$ ms lies well within $[90;140]\;\text{ms}$, the time window of most interest (Sec. 2.3). Although a convergence analysis is completely agnostic to it, the lack of camera saturation along CH#2 at this time is instead crucial for the analyses of Secs. 5.1.3 and 5.1.4. Unless otherwise specified, the results obtained at $t=t_{*}$ apply to any other time instant.
The point $(x_{*},y_{*})$ lies in the volume of the domain, within 3.3 mm of the plasma-facing surface so as to be appreciably affected by a temperature variation (Sec. 4.1.1), and away from any other boundaries, making it suitable for comparison across different cases. This ”local” approach is preferred over integral errors (in the fashion of Sec. 4.3.3) to minimise any near-boundary artifacts which would impair a precise convergence analysis.
Algebra of the energy balance error
By neglecting any discretisation error $o(\Delta\xi)$ with $\xi\in\{x;y;t\}$ in Eqs. (10) and (11), it must follow that:
$\begin{split}\mathcal{E}=&\sum_{i=1}^{N_{x}}\sum_{j=1}^{N_{y}}\varrho c_{\text{p}}(T_{i,j}^{n}-T_{i,j}^{n-1})\times\Delta x\Delta y=\\ =&\sum_{i=1}^{N_{x}}q_{i}^{n}\times\Delta x\Delta t\>.\end{split}$ (17)
Additionally, the errors themselves are such that:
$\begin{split}&o(\Delta\xi)=\sum_{i}A_{i}\>o(\Delta\xi)\;\;\forall A_{i}\in\mathbb{R}\>,\\ &o(\Delta\xi^{0})=o(1)\sim|B|\;\;\forall B\in\mathbb{R}\>,\\ &o(\Delta\xi^{m})\ll o(\Delta\xi^{l})\leq o(1)\;\;\forall m>l\in\mathbb{N}\>,\end{split}$ (18)
Therefore, substituting Eqs. (10) and (11) in Eq. (15), and leveraging Eqs. (17) and (18) yields:
$\begin{split}&\varepsilon_{*}(\mathcal{E})=\frac{\left|\Delta\mathcal{E}_{\text{int*}}-\mathcal{E}_{\text{surf*}}\right|}{\left|\Delta\mathcal{E}_{\text{int*}}+\mathcal{E}_{\text{surf*}}\right|/2}=\\ &=\frac{|\left[o(\Delta x^{2})+o(\Delta y^{2})+o(\Delta t)\right]-\left[o(\Delta x^{2})+o(\Delta y)+o(\Delta t)\right]|}{|\mathcal{E}+\left[o(\Delta x^{2})+o(\Delta y^{2})+o(\Delta t)\right]+\left[o(\Delta x^{2})+o(\Delta y)+o(\Delta t)\right]|}\\ &\sim o(\Delta x^{2})+o(\Delta y)+o(\Delta t)\>.\end{split}$ (19)
Finally, the properties specific to time/space convergence of Sec. Details of time and space convergence, with $\Delta x=\Delta y\Rightarrow o(\Delta x^{2})\ll o(\Delta y)$, lead to the relationships:
time convergence: $\begin{split}&\text{time convergence: $o(\Delta x^{2})\ll o(\Delta t)$}\\ &\Rightarrow\varepsilon_{*}(\mathcal{E})\sim o(\Delta y)+o(\Delta t)\>;\end{split}$ (20)
space convergence: $\begin{split}&\text{space convergence: $o(\Delta t)\ll o(\Delta x^{2})$}\\ &\Rightarrow\varepsilon_{*}(\mathcal{E})\sim o(\Delta y)\>.\end{split}$ (21)